OpenStudy (erod0):

Simplify tan^2theta+1 --------------- sec^2theta

OpenStudy (3mar):

May I help?

OpenStudy (erod0):

\[\frac{ \tan^2\theta+1}{ \sec^2\theta }\]

OpenStudy (erod0):

yes plz

OpenStudy (3mar):

Thank you.

OpenStudy (3mar):

Are you familiar with trigonometric identities?

OpenStudy (3mar):

Can you kindly respond more faster, please?

OpenStudy (3mar):

@erod0

OpenStudy (3mar):

You are funny! The asker asked and walk away! He would take the answer by delivery!

OpenStudy (erod0):

lol wait im sorry

OpenStudy (erod0):

no im not familiar with trigonometric identities

OpenStudy (3mar):

I was waiting for you

OpenStudy (erod0):

i was in the bathroom

OpenStudy (3mar):

Ok No problem.

OpenStudy (3mar):

\[\tan \theta=\frac{ ? }{ ? }\]

OpenStudy (erod0):

i have no idea

OpenStudy (3mar):

\[\tan \theta=\frac{ \sin \theta }{ \cos \theta }\] right?

OpenStudy (3mar):

@erod0

OpenStudy (3mar):

Can you kindly respond more faster, please?

OpenStudy (erod0):

yes

OpenStudy (erod0):

ok im sorry i was peeing

OpenStudy (3mar):

\[(\tan \theta)^2=??\]

OpenStudy (erod0):

is it 1?

OpenStudy (3mar):

Yes it is 1 but you know why or it is just guess?

OpenStudy (erod0):

i remember it from my lecture

OpenStudy (3mar):

So why ""no im not familiar with trigonometric identities""??

OpenStudy (erod0):

well im somewhat familiar with them

OpenStudy (3mar):

You did say that. I thought you are in grade 6 or 8!

OpenStudy (erod0):

im in college lol

OpenStudy (3mar):

So you should know it very very well!!

OpenStudy (erod0):

no not quite. please what is the answer?

OpenStudy (3mar):

The answer is 1 as you and I said!

OpenStudy (3mar):

Got it?

OpenStudy (3mar):

@erod0

OpenStudy (erod0):

yes thank u!

OpenStudy (3mar):

With my pleasure! You are welcome!